OpenStudy (anonymous):

What is the value of the system determinant for the following system of equations? 6x + y = 4 5x + y - 7 = 0 -1 1 31

OpenStudy (anonymous):

i will fan and give someone a medal i always have will someone please help

OpenStudy (anonymous):

6

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

If by system determinant you mean the denominator determinant, then it is: 6 * 1 - 1 * 5 = 6 - 5 = 1

OpenStudy (anonymous):

oh than you so much it helped a ton i didnt understand how to do these problems in general so thanks @mathstudent55

OpenStudy (mathstudent55):

This is how you calculate the value of a 2 by 2 determinant.